Assume you are managing a bank in a country that is holding $34 million in assets and $75 million in liabilities denominated in dollars.

a. Determine the net exposure in dollars.

b. Calculate the gain or loss for each currency below given last year’s spot rate and the current spot rate respectively

    for the U.S. $.

            Euro: 0.78717;    0.9048

            Krone: 5.7538;    6.755

            Columbian Peso: 2,535.253;     3,384.9

            Mexican Peso: 16.9566;      19.3533

            Dinar: .2689;      .30466

Dong: 17,476.5;      23,201.52

c. Convert each of your answers in part b to current dollar equivalents.
Answer #1

a)

Net Exposed Asset = Liabilities - Assets

= $ 75million - $34million

= $41 million

b)

Euro

Gain = 0.9048 - 0.7817 = 0.1231

Krone

Gain = 6.755 - 5.7538 = 1.0012

Columbian Pesso

Gain = 3384.9 - 2535.253 = 849.647

Mexican peso:

Gain = 19.3533 - 16.9566 = 2.3967

Dinar

Gain = 0.30466 - 0.2689 = 0.03576

Dong

Gain = 23201.52 - 17476.5 = 5725.02

c)

Euro Gain in dollars =0.1231*1.16 =$ 0.1428

Krone Gain in dollars =1.0012* 0.1 =$ 0.1

Columbian Pesso Gain in dollars =849.647 *0.00026 =$ 0.22

Mexican peso Gain in dollars =2.3967 *0.045 =$ 0.11

Dinar Gain in dollars =0.03576 * 3.26=$ 0.12

Dong Gain in dollars =5725.02 *0.000043=$ 0.25